OpenStudy (anonymous):

Find the balance in the account 1400 principal earning 6%, compounded semi-annually after 10 years a. 29,680.00 b.2,507.19 c.2,528.56 d.1,468,006,400.00

OpenStudy (anonymous):

@Nurali can you help ??

OpenStudy (nurali):

A=P(1+r/n)^nt Plug in the values: P=1400, r=6/100=0.6, n=2 and t=10 A=1400(1+.06/2)^2(10) A=$2528.56

OpenStudy (anonymous):

thank you ^_^